# Cross-correlation scheme for quantum optical coherence tomography based on Michelson interferometer

### v1

@@ I Introduction
The resolution of far-field imaging systems is limited by the optical radiation wavelength, and the numerical aperture (NA). While the transverse resolution is inversely proportional to NA, the axial resolution is inversely proportional to the square of NA [1]. This limitation is particularly critical in applications, where NA is significantly limited, such as retina imaging through an eye pupil [2].
@@
The technique that addresses this limitation is Optical coherence tomography (OCT), introduced in Ref. [3] and widely described in Ref. [2]. In the time-domain OCT scheme, presented in Fig. 1a, the sample is placed in one arm of a Michelson interferometer (MI), while the other arm’s delay, $\tau$, is varied. Each reflective layer of the sample generates an interferogram $M_{\text{OCT}}(\tau)$, allowing reconstruct the structure of the sample – coordinates and reflectivities of the layers. The interferogram, obtained from a single layer can be considered as a point spread function (PSF) of the imaging system. Its width determines the axial resolution, independent of NA but limited by the coherence length of the optical source. Thus, a low-coherent broadband source is required for fine axial resolution. However, any dispersion in the sample or setup components broadens the interferogram, severely limiting the scanning depth of OCT unless specific sample dispersion properties are known and compensated.
@@
Quantum optical coherence tomography (QOCT) [4], based on Hong-Ou-Mandel (HOM) interference [5], presented in Fig. 1b, overcomes the dispersion broadening issue and also provides double resolution. In QOCT, each reflective layer produces a dip $M_{\text{HOM}}(\tau)$, where the width, inversely proportional to the bandwidth of the spontaneous parametric down-conversion (SPDC) [6] field, defines the system’s axial resolution. Notably, QOCT is insensitive to group velocity dispersion if the SPDC pump field is sufficiently narrowband [7], and it provides twice the axial resolution of OCT with the same source spectrum. Additionally, as mentioned in Ref. [5], HOM interference does not require any phase stabilization or sub-wavelength scanning steps, and the interferogram pattern $M_{\text{HOM}}(\tau)$ is much clearer and easier to process compared to $M_{\text{OCT}}(\tau)$.
@@
Despite its advantages, the practical application of QOCT remains limited due to the long measurement times required by the low brightness of broadband SPDC sources. Bright sources are more effortlessly produced via collinear phase matching, which allows for effective use of longer crystal lengths. As mentioned in Ref [8], for broadband sources, degenerate phase matching types-I or 0 are preferable, though this leads to both photons being in the same optical mode, preventing their deterministic separation for HOM interference.
@@
Recent studies suggest that QOCT can also be implemented using a brighter, broader, and more adjustable collinear type-I/type-0 SPDC source coupled with a MI with autocorrelation measurement scheme [9, 10, 11, 12, 13]. In this configuration (Fig. 1c), the interferometer’s output signal is split, and the coincidence count rate interferogram $M_{\text{MI}}^{\text{(auto)}}(\tau)$, is measured.
@@
However, achieving sub-micrometer axial resolution [13] requires an ultra-broadband SPDC source, which causes overlap between the target and parasitic components in Fourier domain. Phase instabilities further broaden these features, making post-processing more challenging, reducing the usable spectral bandwidth, and ultimately limiting the axial resolution.
To overcome these limitations, we explore and experimentally demonstrate a cross-correlation MI-based QOCT scheme (Fig. 1d). This scheme combines dispersion cancellation, doubled resolution, and robustness to phase instability of the standard HOM approach with the high brightness of collinear SPDC, while eliminating overlap between parasitic Fourier components of the interferogram. By monitoring coincidences between the MI’s two output ports (one of which is recycled to the input via a circulator), we obtain the PSF interferogram $M_{\mathrm{MI}}^{\mathrm{(cross)}}(\tau)$. Its Fourier-transform $\tilde{M}_{\mathrm{MI}}^{\mathrm{(cross)}}(\omega)$ contains only a narrow pump-frequency peak $\tilde{M}_{2}$ and the target HOM term $\tilde{M}_{0}$. Because the parasitic peak is well separated from the HOM dip, it can be removed by simple Fourier filtering or by averaging over high-frequency oscillations during a coarse mechanical scan – without limiting SPDC bandwidth or requiring sub-wavelength path control.
This approach paves the way for practical QOCT applications. While a similar scheme was recently applied to refractometry [14], here we apply cross-correlation QOCT to defectoscopy.
The paper is organized as follows. Section II provides a full theoretical description of both autocorrelation and cross-correlation MI-based QOCT schemes and highlights the advantages of the latter. Section III describes the experimental setup, combining OCT and both QOCT schemes. Section IV compares resolution and signal-to-noise ratio across all three techniques under various conditions, and applies our cross-correlation QOCT to resolve a $\sim$110 µm air gap between highly dispersive silicon windows – far below the classical OCT resolution.
@@ II Theory
Figure 2: Michelson Interferometer.
In this section we exploit the approach [13] to describe quantum interference relating the MI-based QOCT schemes Fig. 1 c and d and theoretically compare them.
Consider an interference of a photon pair in MI. The interferometer scheme is illustrated in Fig. 2. It consists of a 50/50 beam splitter, a reflective sample, and a translated mirror in a reference channel. It connects four input and four output optical modes. Photon creation operators in these modes are denoted as $\hat{a}^{\dagger}_{\text{in}}$, $\hat{a}^{\dagger}_{\text{out}}$, $\hat{b}^{\dagger}_{\text{in}}$, $\hat{b}^{\dagger}_{\text{out}}$, $\hat{c}^{\dagger}_{\text{in}}$, $\hat{c}^{\dagger}_{\text{out}}$, $\hat{d}^{\dagger}_{\text{in}}$, $\hat{d}^{\dagger}_{\text{out}}$. The beam splitter connects these operators as follows:
@@
Denote the sample response function as $H(\omega)$ and delay in the reference channel as $\tau$. Then input and output modes are connected as follows:
@@
The general description of the photon pair MI-interference is presented in A. Here we consider two assumptions.
  • The sample has a single layer without dispersion with the following response function:
    $H\left(\omega\right)=r{{\operatorname{e}}^{\operatorname{i}\omega T}},\quad{{r}^{2}}\equiv R-\text{layer reflectivity}.$ (3)
    Here $\omega T=kd=\frac{n\omega d}{c}$, where $d$ is the sample thickness, $n$ is the refractive index and $c$ is the speed of light. So $T=\frac{nd}{c}$.
  • The input state is a collinear degenerate SPDC state with a type-0/type-I phase matching:
    $\left|{{\Psi}_{\text{in}}}\right\rangle=\frac{1}{\sqrt{2}}\iint{\text{d}{{\omega}_{s}}\text{d}{{\omega}_{i}}}f\left({{\omega}_{s}},{{\omega}_{i}}\right)\hat{a}_{\text{in}}^{\dagger}\left({{\omega}_{s}}\right)\hat{a}_{\text{in}}^{\dagger}\left({{\omega}_{i}}\right)\left|\text{vac}\right\rangle,$ (4)
    where
    ${\left|f({{\omega}_{s}},{{\omega}_{i}})\right|}^{2}=\underbrace{G(\omega_{1}-\omega_{2}|0,\Delta)}_{\text{phase matching}}\times\underbrace{G(\omega_{1}+\omega_{2}|\omega_{\text{p}},\delta)}_{\text{pump}}$ (5)
    is a Gaussian approximation of a photon pair spectral amplitude [15, 16], visualized in Fig. 3. Here
    $G(\omega|\omega_{\text{c}},\sigma)\equiv\frac{1}{\sqrt{\pi}\sigma}\exp\left[\frac{(\omega-\omega_{\text{c}})^{2}}{2\sigma^{2}}\right]$ (6)
    is a normalized Gaussian distribution, $\delta$ represents the pump bandwidth, and $\Delta$ is the phase matching bandwidth (in this section we define the bandwidth as a standard deviation of the Gaussian distribution).
Figure 3: Two-photon spectral distribution |$\absolutevalue{f(\omega_{\text{s}},\omega_{\text{i}})}$ | for degenerate case.
Traditional MI-based autocorrelation QOCT scheme uses the detector realizing projection on the two-photon state in the mode $b$:
@@
As an alternative, we consider a cross-correlation scheme based on the detector realizing projection on the state with two photons in different modes $a$ and $b$:
@@
The probability to register one photon in the mode $m$ and another one in the mode $n$, depending on the delay in the reference channel $\tau$ (two-photon interferogram) is given by:
@@
Substituting Eqs. (17) into Eq. (8) we obtain the final interferograms for autocorrelation $(bb)$ and cross-correlation $(ab)$ measurement scheme:
$M_{\text{MI}}^{\text{(auto)}}$ $\equiv M^{(bb)}=\frac{M_{c}+M_{0}-M_{1}+M_{2}}{16},$ (9a)
$M_{\text{MI}}^{\text{(cross)}}$ $\equiv M^{(ab)}=\frac{M_{c}-M_{0}-M_{2}}{8}.$ (9b)
@@
${{M}_{c}}=\left(1+R\right)^{2}$ (10a)
is a constant term,
${{M}_{0}}(\tau)=2R\exp\left[-\frac{{{\left(T-\tau\right)}^{2}}}{2{{\Delta}^{-2}}}\right]$ (10b)
is a term, related to HOM interference, presented in Fig. 1b,
${{M}_{1}}(\tau)=4r\left(1+R\right)\exp\left[-\frac{{{\left(T-\tau\right)}^{2}}}{8\Delta_{+}^{-2}}\right]\cos\left[\frac{{{\omega}_{\text{p}}}\left(T-\tau\right)}{2}\right]$ (10c)
with $\Delta_{+}^{2}\equiv\delta^{2}+\Delta^{2}\approx\Delta^{2}$ is a term related to a single-photon Michelson interference, shown in Fig. 1a, and
${{M}_{2}}(\tau)=2R\ \exp\left[-\frac{{{\left(T-\tau\right)}^{2}}}{2{{\delta}^{-2}}}\right]\cos\left[{{\omega}_{\text{p}}}\left(T-\tau\right)\right]$ (10d)
@@
|$\absolutevalue{\tilde{M}_{0}(\omega)}$ | $=2\sqrt{\pi}R\times G(\omega|0,\Delta),$ (11a)
|$\absolutevalue{\tilde{M}_{1}(\omega)}$ | $=2\sqrt{\pi}r(1+R)\times G({\omega|\omega_{\text{p}}}/{2},{\Delta_{+}}/{2}),$ (11b)
|$\absolutevalue{\tilde{M}_{2}(\omega)}$ | $=\sqrt{\pi}R\times G(\omega,\omega_{\text{p}},\delta).$ (11c)
The target $M_{0}$ term can be easily extracted, if it does not overlap with $M_{1}$ term in Fourier domain, which requires $\Delta\ll\omega_{\text{p}}/3$. However, to achieve fine axial resolution ${\rm d}z\sim c/\Delta$ one needs to use a broadband SPDC source. For example, in Ref. [13] authors used the autocorrelation MI-QOCT scheme with an SPDC source, covering the range 684–992 nm, which corresponds to $\Delta\approx 0.08\omega_{\text{p}}$. It resulted in non-negligible overlap between $\tilde{M}_{0}$ and $\tilde{M}_{1}$ peaks, which disimproved the total QOCT resolution and signal-to-noise ratio.
However, there are no $M_{1}$ term in the cross-correlation interferogram $M_{\text{MI}}^{\text{(cross)}}$. It has only one parasitic term $M_{2}$ with a spectrally separated peak and a narrow bandwidth $\delta$, which can be easily eliminated ether by a Fourier-processing of the interferogram, ether by an averaging over the delay ${\rm d}\tau\sim 2\pi/\omega_{\text{p}}$ during a continuous delay scan.
Additionally, as follows from Eq. (9), the target $M_{0}$ term in the interferogram $M_{\text{MI}}^{\text{(cross)}}$ has a twice bigger amplitude in comparison with $M_{\text{MI}}^{\text{(auto)}}$ and a negative sign, corresponding to the dip shape. The $M_{1}$ term in the interferogram $M_{\text{MI}}^{\text{(auto)}}$ has a positive sign, corresponding to a peak shape, since it is equivalent to the probability to detect two photons in one output in the standard HOM-interference. Moreover, since usually two-photon probability is measured with a beam splitter and two single-photon detectors, as shown in Fig. 1c, it leads to addition 50% loss in coincidence rate, thus the measured $M_{1}$ amplitude in the cross-correlation scheme is 4 times bigger than in the autocorrelation scheme.

### v2

@@ I Introduction
The resolution of far-field imaging systems is limited by the optical radiation wavelength and the numerical aperture (NA). While the transverse resolution is inversely proportional to NA, the axial resolution is inversely proportional to the square of NA [pawley_handbook_2006]. This limitation is particularly critical in applications where NA is significantly limited, such as retinal imaging through an eye pupil [drexler_optical_2015].
@@
The technique that addresses this limitation is optical coherence tomography (OCT), introduced in Ref. [huang_optical_1991] and widely described in Ref. [drexler_optical_2015]. In the time-domain OCT scheme, presented in Fig. 1a, the sample is placed in one arm of a Michelson interferometer (MI), while the other arm’s delay, $\tau$, is varied. Each reflective layer of the sample generates an interferogram $M_{\text{OCT}}(\tau)$, allowing one to reconstruct the structure of the sample – coordinates and reflectivities of the layers. The interferogram obtained from a single layer can be considered as a point-spread function (PSF) of the imaging system. Its width determines the axial resolution, independent of NA but limited by the coherence length of the optical source. Thus, a low-coherence broadband source is required for fine axial resolution. However, any dispersion in the sample or setup components broadens the interferogram, severely limiting the scanning depth of OCT unless specific sample dispersion properties are known and compensated.
Several dispersion‐cancellation strategies for OCT have been demonstrated. Some – like the hardware phase‐compensation method [hitzenberger_dispersion_1999] – require precise a priori knowledge of the sample’s dispersion and only cancel dispersion at a single imaging depth. Others rely on elaborate nonlinear‐optics implementations, for example phase‐conjugation or chirped‐pulse schemes [erkmen_phaseconjugate_2006, kaltenbaek_quantuminspired_2008, legouet_experimental_2010, lavoie_quantumoptical_2009], which demand high‐power lasers and complex setups. Other dispersion-cancellation approaches rely on intensive numerical spectral post-processing [banaszek_blind_2007, resch_classical_2007, ryczkowski_experimental_2016, shirai_intensityinterferometric_2014, jensen_alldepth_2018, liu_dispersion_2020], which makes them intrinsically sensitive to interferometric phase and timing instabilities.
@@
Quantum optical coherence tomography (QOCT) [abouraddy_quantum-optical_2002, Nasr2003], based on Hong-Ou-Mandel (HOM) interference [Hong1987b], presented in Fig. 1b, overcomes the dispersion broadening issue and also provides double resolution. In QOCT, each reflective layer produces a dip $M_{\text{HOM}}(\tau)$, where the width, inversely proportional to the bandwidth of the spontaneous parametric down-conversion (SPDC) [Klyshko1988a] field, defines the system’s axial resolution. Notably, QOCT is insensitive to group-velocity dispersion if the SPDC pump field is sufficiently narrowband [okanoDispersionCancellationHighresolution2013], and it provides twice the axial resolution of OCT with the same source spectrum. Additionally, as mentioned in Ref. [Hong1987b], HOM interference does not require any phase stabilization or sub-wavelength scanning steps, and the interferogram pattern $M_{\text{HOM}}(\tau)$ is much clearer and easier to process compared to $M_{\text{OCT}}(\tau)$.
@@
Despite its advantages, the practical application of QOCT remains limited due to the long measurement times required by the low brightness of broadband SPDC sources. Bright sources are more effortlessly produced via collinear phase matching, which allows for effective use of longer crystal lengths. As mentioned in Ref. [Katamadze2022], for broadband sources, degenerate phase matching types-I or 0 are preferable, though this leads to both photons being in the same optical mode, preventing their deterministic separation for HOM interference.
@@
Recent studies suggest that QOCT can also be implemented using a brighter, broader [baek_SpectralProperties2008], and more adjustable collinear type-0/type-I SPDC source coupled to an MI with an autocorrelation measurement scheme [odateTwophotonQuantumInterference2005, lopez-mago_coherence_2012, lopez-mago_quantum-optical_2012, yoshizawa_telecom-band_2014, katamadze_broadband_2025]. In this configuration (Fig. 1c), the interferometer’s output signal is split, and the coincidence count rate interferogram $M_{\text{MI}}^{\text{(auto)}}(\tau)$, is measured.
@@
However, achieving sub-micrometer axial resolution [katamadze_broadband_2025] requires an ultra-broadband SPDC source, which causes overlap between the target and parasitic components in the Fourier domain. Phase instabilities further broaden these features, making post-processing more challenging, reducing the usable spectral bandwidth, and ultimately limiting the axial resolution.
To overcome these limitations, we explore and experimentally demonstrate a cross-correlation MI-based QOCT scheme (Fig. 1d). This scheme combines dispersion cancellation, doubled resolution, and robustness to phase instability of the standard HOM approach with the high brightness of collinear SPDC, while eliminating overlap between parasitic Fourier components of the interferogram. By monitoring coincidences between the MI’s two output ports (one of which is recycled to the input via a circulator), we obtain the PSF interferogram $M_{\mathrm{MI}}^{\mathrm{(cross)}}(\tau)$. Its Fourier transform $\tilde{M}_{\mathrm{MI}}^{\mathrm{(cross)}}(\omega)$ contains only a narrow pump-frequency peak $\tilde{M}_{2}$ and the target HOM term $\tilde{M}_{0}$. Because the parasitic peak is well separated from the HOM dip, it can be removed by simple Fourier filtering or by averaging over high-frequency oscillations during a coarse mechanical scan – without limiting SPDC bandwidth or requiring sub-wavelength path control.
The cross-correlation approach was recently demonstrated for high-precision refractive-index measurements in optical fibers [reisner_quantumlimited_2022], however, in that application the PSF width is not critical because each scan measures a single path length and the peak can be located with sub-PSF precision. By contrast, here we provide a comprehensive study of a Michelson-based cross-correlation QOCT: we compare its resolution and signal-to-noise ratio (SNR) with both autocorrelation QOCT and classical OCT, analyze how the axial resolution depends on the source spectrum, demonstrate enhanced robustness to phase noise and optical dispersion, and finally apply the method to defectoscopy – a task where PSF narrowing and dispersion cancellation are essential.
The paper is organized as follows. Section II presents the theoretical description of classical OCT and of both autocorrelation and cross-correlation MI-based QOCT schemes and highlights the advantages of the latter. Section III describes the experimental setup that implements both QOCT schemes alongside a classical OCT reference. Section IV compares resolution and SNR across the three techniques and includes a defectoscopy demonstration.
@@ II Theory
[…]
In this section we follow the approach of Ref. [katamadze_broadband_2025] to describe quantum interference in the MI-based QOCT schemes of Fig. 1a, c, and d and to compare them theoretically.
Consider interference of a photon pair in MI. The interferometer scheme is illustrated in Fig. 2. It consists of a 50/50 beam splitter, a reflective sample, and a translated mirror in a reference channel. It connects four input and four output optical modes. Photon creation operators in these modes are denoted as $\hat{a}^{\dagger}_{\text{in}}$, $\hat{a}^{\dagger}_{\text{out}}$, $\hat{b}^{\dagger}_{\text{in}}$, $\hat{b}^{\dagger}_{\text{out}}$, $\hat{c}^{\dagger}_{\text{in}}$, $\hat{c}^{\dagger}_{\text{out}}$, $\hat{d}^{\dagger}_{\text{in}}$, $\hat{d}^{\dagger}_{\text{out}}$. The beam splitter connects these operators as follows:
@@
Let $H(\omega)$ denote the sample response function and $\tau$ the delay in the reference channel. Then input and output modes are connected as follows:
@@
The general description of the single-photon and photon pair MI-interference is presented in Appendix A. Here we make two assumptions.
  • The sample has a single layer without dispersion, with the response function
    $H(\omega)=r\,\operatorname{e}^{\operatorname{i}\omega T},\quad r^{2}\equiv R,$ (3)
    where $R$ is the layer reflectivity. Here, $\omega T=kd=\tfrac{n\omega d}{c}$, with $d$ the layer thickness, $n$ the refractive index, and $c$ the speed of light. Thus, $T=\tfrac{nd}{c}$.
  • The input state consists of photons with the same polarization, propagating in input mode $a$ and having Gaussian spectral distribution.
First, we consider a classical or single-photon state with a Gaussian spectral distribution
$G(\omega|\omega_{0},\sigma)\equiv\frac{1}{\sqrt{2\pi}\sigma}\exp\left[-\frac{(\omega-\omega_{0})^{2}}{2\sigma^{2}}\right]$ (4)
with the central frequency $\omega_{0}$ and the bandwidth $\sigma$ (in this section we define the bandwidth as a standard deviation of the Gaussian distribution). Measuring the output intensity (count rate) in the mode $b$ as a function of $\tau$ yields the following interferogram:
$M^{(b)}(\tau)=\frac{1+R}{4}-\frac{r}{2}\exp\left[-\frac{{{\left(T-\tau\right)}^{2}}}{2\sigma^{-2}}\right]\cos\left[{{{\omega}_{0}}\left(T-\tau\right)}\right],$ (5)
presented in Fig. 1a (see Appendix A for details). This interferogram represents the PSF of classical OCT.
For QOCT we consider a collinear degenerate SPDC state with type-0/type-I phase matching:
$\left|{{\Psi}_{\text{in}}}\right\rangle=\frac{1}{\sqrt{2}}\iint{\text{d}{{\omega}_{\text{s}}}\text{d}{{\omega}_{\text{i}}}}F\left({{\omega}_{\text{s}}},{{\omega}_{\text{i}}}\right)\hat{a}_{\text{in}}^{\dagger}\left({{\omega}_{\text{s}}}\right)\hat{a}_{\text{in}}^{\dagger}\left({{\omega}_{\text{i}}}\right)\left|\text{vac}\right\rangle,$ (6)
where ${F}({{\omega}_{\text{s}}},{{\omega}_{\text{i}}})$ is a spectral amplitude and
${\left|F({{\omega}_{\text{s}}},{{\omega}_{\text{i}}})\right|}^{2}=2\underbrace{G(\omega_{\text{s}}-\omega_{\text{i}}|0,\Delta)}_{\text{phase matching}}\times\underbrace{G(\omega_{\text{s}}+\omega_{\text{i}}|\omega_{\text{p}},\delta)}_{\text{pump}}$ (7)
is a Gaussian approximation of the photon-pair spectral distribution [Mikhailova2008, Fedorov2009], visualized in Fig. 3. Here $\Delta$ represents the phase matching bandwidth and $\delta\ll\Delta$ is the pump bandwidth.
The corresponding marginal single-photon spectral distribution also has a Gaussian form:
$\int{\text{d}{{\omega}_{\text{i}}}{\left|F({{\omega}_{\text{s}}},{{\omega}_{\text{i}}})\right|}^{2}}={G(\omega_{\text{s}}|\omega_{\text{p}}/2,\Delta_{+}/2)},$ (8)
with the bandwidth twice narrower than $\Delta_{+}\equiv\sqrt{\Delta^{2}+\delta^{2}}\approx\Delta$.
Figure 3: Two-photon spectral distribution $|\absolutevalue{f(\omega_{\text{s}},\omega_{\text{i}})}$ for degenerate case.
The traditional MI-based autocorrelation QOCT scheme uses a detector that realizes a projection onto the two-photon state in mode $b$:
@@
As an alternative, we consider a cross-correlation scheme based on a detector that projects onto the state with two photons in different modes $a$ and $b$:
@@
The probability to register one photon in the mode $m$ and another one in the mode $n$, as a function of the delay $\tau$ in the reference channel (the two-photon interferogram) is given by:
@@
Substituting Eqs. (19) into Eq. (10) we obtain the final interferograms for autocorrelation $(bb)$ and cross-correlation $(ab)$ measurement schemes:
$M_{\text{MI}}^{\text{(auto)}}$ $\equiv M^{(bb)}=\frac{M_{c}+M_{0}-M_{1}+M_{2}}{4},$ (11a)
$M_{\text{MI}}^{\text{(cross)}}$ $\equiv M^{(ab)}=\frac{M_{c}-M_{0}-M_{2}}{2}.$ (11b)
@@
${{M}_{c}}=\frac{1}{4}\left(1+R\right)^{2}$ (12a)
is a constant term. The term
${{M}_{0}}(\tau)=\frac{1}{2}R\exp\left[-\frac{{{\left(T-\tau\right)}^{2}}}{2{{\Delta}^{-2}}}\right]$ (12b)
is related to HOM interference, presented in Fig. 1b.
The next term
${{M}_{1}}(\tau)=r\left(1+R\right)\exp\left[-\frac{{{\left(T-\tau\right)}^{2}}}{8\Delta_{+}^{-2}}\right]\cos\left[\frac{{{\omega}_{\text{p}}}\left(T-\tau\right)}{2}\right]$ (12c)
is a term related to a single-photon Michelson interference. It fits up to a constant factor with the second term of the single-photon interferogram (5) for the single-photon Gaussian spectral distribution (8). The last term
${{M}_{2}}(\tau)=\frac{1}{2}R\ \exp\left[-\frac{{{\left(T-\tau\right)}^{2}}}{2{{\delta}^{-2}}}\right]\cos\left[{{\omega}_{\text{p}}}\left(T-\tau\right)\right]$ (12d)
@@
$|\displaystyle\absolutevalue{\tilde{M}_{0}(\omega)}$ $=\frac{\sqrt{\pi}R}{2}\ G(\omega|0,\Delta),$ (13a)
$|\displaystyle\absolutevalue{\tilde{M}_{1}(\omega)}$ $=\frac{\sqrt{\pi}r(1+R)}{2}\ G\!\left(\omega\left|\frac{\omega_{\text{p}}}{2}\right.,\frac{\Delta_{+}}{2}\right),$ (13b)
$|\displaystyle\absolutevalue{\tilde{M}_{2}(\omega)}$ $=\frac{\sqrt{\pi}R}{4}\ G(\omega,\omega_{\text{p}},\delta).$ (13c)
The target $M_{0}$ term can be easily extracted if it does not overlap with $M_{1}$ term in the Fourier domain, which requires $\Delta\ll\omega_{\text{p}}/3$. However, to achieve fine axial resolution ${\rm d}z\sim c/\Delta$ one needs to use a broadband SPDC source. For example, in Ref. [katamadze_broadband_2025] the autocorrelation MI-QOCT scheme with an SPDC source was used, covering the range 684–992 nm, which corresponds to $\Delta\approx 0.08\,\omega_{\text{p}}$. It resulted in non-negligible overlap between $\tilde{M}_{0}$ and $\tilde{M}_{1}$ peaks, which degraded the total QOCT resolution and SNR.
However, there is no $M_{1}$ term in the cross-correlation interferogram $M_{\text{MI}}^{\text{(cross)}}$. It has only one parasitic term $M_{2}$ with a spectrally separated peak and a narrow bandwidth $\delta$, which can be easily eliminated either by a Fourier-processing of the interferogram, or by averaging over the delay ${\rm d}\tau\sim 2\pi/\omega_{\text{p}}$ during a continuous delay scan.
Additionally, as follows from Eq. (11), the target $M_{0}$ term in the interferogram $M_{\text{MI}}^{\text{(cross)}}$ has a twice the amplitude in comparison with $M_{\text{MI}}^{\text{(auto)}}$ and a negative sign, corresponding to the dip shape. The $M_{1}$ term in the interferogram $M_{\text{MI}}^{\text{(auto)}}$ has a positive sign, corresponding to a peak shape, since it is equivalent to the probability to detect two photons in one output in the standard HOM-interference. Moreover, since usually two-photon probability is measured with a beam splitter and two single-photon detectors, as shown in Fig. 1c, it leads to an additional 50% loss in coincidence rate, thus the measured $M_{1}$ amplitude in the cross-correlation scheme is 4 times bigger than in the autocorrelation scheme.